OpenStudy (anonymous):

PLEASE HELP what are the zeros of the function? y=(x-3(x+2)(x-2)

OpenStudy (anonymous):

set \(x-3=0\) and solve also solve \(x+2=0\) and \(x-2=0\) you can do it in your head

OpenStudy (anonymous):

thanks